OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line. (2,-8); 4x-7y=6

OpenStudy (anonymous):

\[y=\frac{4}{7}x-\frac{64}{7}\]

OpenStudy (asadkarim7):

parallel lines same slope i.e 7/ 4 put it into y=mx+c -8 = 7/4 ( 2) +c c= - 11 1/2 y = 7/4 x - 23/2

OpenStudy (anonymous):

parallel lines have the same slope asadkarim7. Perpendicular lines have inverse slopes

OpenStudy (anonymous):

thank you both of you! beehive that was the correct answer, you saved my life. lol